A loan of $50,000 is paid off by paying 6% reducible interest. Annual repayments of $6,000 are paid at the end of the year.
a. Determine the amount that is still owing at the end of the third year.
b. How much interest is paid altogether?
Thank you :)

Answer by Theo(13342)

the loan is 50,000.
the interest rate is 6% per year on the remaining balance of the loan.
the payment is at the end of each year.

here's what happens.

at the beginning of the first year, the remaining balance is 50,000

at the end of the first year ...
interest is 50,000 * .06 = 3,000
payment is 6,000
remaining balance is 50,000 + 3,000 - 6,000 = 47,000

at the end of the second year ...
interest is 47,000 * .06 = 2,820
payment is 6,000
remaining balance is 47,000 + 2,820 - 6,000 = 43,820

at the end of the third year ...
interest is 43,820 * .06 = 2,629.2
payment is 6,000
remaining balance is 43,820 + 2,629.2 - 6,000 = 40,449.2

the total interest paid over the 3 year period is 3,000 + 2,820 + 2,629.2.
this is equal to 8,449.2